# A Deterministic Information Bottleneck Method for Clustering Mixed-Type Data

### v5

@@ 1 Introduction
The literature on clustering offers several taxonomies of algorithms, categorising them based on key features such as the type of output (e.g., hard or fuzzy clustering) or the methodological approach (e.g., model-based or centroid-based). For mixed-type data, model-based approaches have been widely explored, with the Gaussian-Multinomial mixture model [4] standing out as a prominent example. Latent variable methods, which introduce unobserved variables to capture underlying patterns, have also been effectively applied to clustering mixed-type data [5]. Non-model-based approaches, often referred to as distance-based methods, provide another robust framework for clustering heterogeneous data. Among these, the K-Prototypes algorithm [6] is particularly notable for combining the strengths of K-Means clustering for continuous variables with the K-Modes method for categorical variables. Ji et al. [7] further enhanced this approach by incorporating attribute weights in the K-Prototypes algorithm through entropy-based feature weighting, demonstrating improved performance particularly on data sets with varying feature importance.
Building on this foundation, recent work by Rezaei and Daneshpour [8] proposed a three-fold approach in mixed-data clustering: first, they introduced a density-based method for intelligent selection of initial cluster centers, reducing the need for random initialization; second, they developed a more nuanced similarity measure that considers the ratio of similar features between observations, where similarity is determined through a user-defined threshold rather than strict matching; and third, they proposed a clustering algorithm that combines both distance metrics and feature similarity counts when assigning observations to clusters, particularly when distances to cluster centers are comparable. Their approach demonstrated improved clustering accuracy especially for data sets with diverse feature types, showing particular strength in handling smaller, feature-rich mixed data sets. However, this method shows limitations when dealing with large-scale data sets with few classes, and its performance heavily depends on careful tuning of the threshold parameter, making it less robust for general-purpose applications. More recently, Kar et al. [9] proposed EDMD (Entropy-based Dissimilarity Measure for Mixed-categorical Data), which employs Boltzmann’s entropy definition rather than Shannon entropy to better capture the inherent disorder in categorical data. Their method separately treats nominal and ordinal attributes by computing microstates - the number of possible arrangements between categorical values - to determine dissimilarity. They also incorporate attribute weighting based on the statistical significance of different features. While EDMD demonstrated strong performance particularly on data sets with diverse feature types and ordinal features, its effectiveness requires data sets with sufficiently high feature diversity. Further recent developments in distance-based approaches include the work of Mousavi and Sehhati [10], who proposed GUDMM (Generalized Unified Distance Metric for Mixed-type data), which systematically incorporates both intra-attribute and inter-attribute relationships. Their method utilizes Jensen-Shannon divergence and mutual information to capture dependencies between different attribute types, while also distinguishing between nominal and ordinal features. For numerical attributes, they introduced a modified Mahalanobis distance that considers attribute associations. While GUDMM demonstrated strong performance across various dataset types, particularly when attributes exhibit significant dependencies, its reliance on pairwise distance calculations can incur high computational costs for high-dimensional data.
@@
The remainder of this paper is structured as follows: Section 2 presents the Deterministic Information Bottleneck method tailored for mixed-type data (DIBmix), detailing its theoretical framework and algorithmic implementation. Section 3 outlines the selection process of hyperparameter values and Section 4 discusses the simulations performed on artificial data to benchmark the proposed method against other established clustering techniques. In Section 5, we apply the DIBmix method to publicly available data sets and analyse its performance. Finally, Section 6 summarises our findings and suggests avenues for future research.
2 Methodology
@@
$q^{*}(t\mid x)=\operatornamewithlimits{argmin}\limits\limits_{q(t\mid x)}H(T)-\beta I(Y;T),$ (1)
@@
$\quad\times\prod\limits_{j=p_{c}+p_{u}+1}^{p}K_{o}\left(x_{q,j}=x_{r,j};\nu_{j}\right),$
$\quad q,r=1,\ldots,n.$
@@
31exCompute perturbed similarity matrix $\mathbf{P^{\prime}}=\mathbf{P_{Y\mid X}}$ using generalised product kernels with the bandwidths $\boldsymbol{\theta}=(\boldsymbol{s}^{\intercal},\boldsymbol{\lambda}^{\intercal},\boldsymbol{\nu}^{\intercal})^{\intercal}$.
@@
8 while $m\leq m^{\text{max}}$ and $\mathrm{converged}\neq 1$ do
@@
10
11  1exif $\mathbf{Q^{(m)}_{T\mid X}}=\mathbf{Q^{(m-1)}_{T\mid X}}$ then
@@
12
13     1exbreak
14 $\mathbf{q^{(m)}_{T}}\leftarrow 1/n\times\mathbf{Q^{(m)}_{T\mid X}}\boldsymbol{1}_{n}$
@@ 3.1 Bandwidth Selection
For continuous variables, we determine a reasonable bandwidth value $s$ (assuming $\boldsymbol{s}=s(1,\ldots,1)^{\intercal}$) by seeking a high degree of local smoothing. This is quantified using the *average nearest-neighbour kernel ratio* $\zeta(\boldsymbol{s})$, defined similarly to Expression (7):
@@ 4 Simulations on Artificial Data
Data generation was performed using the MixSim function from the MixSim package [34], with each scenario replicated 50 times. For continuous data, cluster overlap is defined as in [35]. To incorporate nominal features, quartile-based discretisation was applied. In total, 28,800 data sets were generated. Cluster recovery was assessed using the Adjusted Rand Index (ARI) [36] and the Adjusted Mutual Information (AMI) [37]. We only report the ARI values here for brevity; the same conclusions are drawn by considering the AMI values, which are available in the GitHub link provided earlier.
Figure 2: Violin/box plots of Adjusted Rand Index values by method.
@@
Figure 2 illustrates the distribution of ARI values across the five methods. The median ARI is highest for DIBmix, followed by KAMILA, K-Prototypes, and FAMD with K-Means, suggesting that these methods are more consistent in producing high-quality cluster partitions. Conversely, PAM with Gower’s dissimilarity exhibits a wider spread of ARI values, including several low outliers, indicating less stable performance. A repeated measures analysis of variance in which the four competing algorithms (KAMILA, K-Prototypes, FAMD with K-Means, and PAM/Gower) are compared individually with DIBmix also verifies the improvement in cluster recovery performance with the latter. More precisely, comparing the average performance of DIBmix with that of the competing methods in terms of ARI and AMI scores yielded effect sizes that all correspond to partial $\eta^{2}$ values of over 0.14, the latter value being a rule-of-thumb for interpreting large effects [38]. The partial $\eta^{2}$ values are summarised in Table 1; $p$-values are not reported as these are almost zero due to the large number of comparisons.
@@
Figure 3 provides a detailed comparison of mean ARI values across varying experimental conditions. Overall, DIBmix outperforms the other methods in a majority of scenarios. However, its performance is relatively less competitive when the proportion of categorical variables is high. This may be attributable to the bandwidth selection strategy outlined in subsection 3.1; alternative bandwidth choices could potentially enhance performance in these cases. Despite this limitation, DIBmix still performs comparably to other methods in high-categorical scenarios, where most methods, except FAMD, struggle. The dimensionality reduction step in FAMD likely provides an advantage in handling data sets dominated by categorical features. Another noteworthy observation is DIBmix’s robustness to cluster imbalances. Unlike KAMILA, K-Prototypes, and FAMD, DIBmix effectively handles data sets with unbalanced cluster sizes. This robustness arises from the entropy term in its objective function, which minimises for imbalanced clusters, thus allowing exploration of diverse partition structures.
@@ 5 Applications to Publicly Available Data
While artificial data provides controlled settings, it does not cover the full range of scenarios encountered in practice. We therefore complement those experiments with evaluations on publicly available data sets from the UCI Machine Learning Repository [39]. These data sets were originally designed for classification tasks, providing a benchmark for assessing clustering accuracy. The ARI values, averaged over one hundred random initialisations, are summarised in Table 2, with the best result for each dataset shown in bold.
@@
Adult/Census Income 0.1739 0.1514 0.0864 0.1558 0.1007
@@
    Best performance occurs when data sets contain a balanced mix of categorical and continuous variables. For instance, in the Heart Disease, Adult, Statlog, and Australian Institute of Sport data sets, which all feature a substantial but not overwhelming proportion of categorical variables, DIBmix outperformed its competitors.
@@
    Performance decreases at the extremes. When data sets are dominated by categorical variables (e.g., Dermatology, with nearly 97% categorical features), or predominantly continuous (e.g., Australian Institute of Sport, Echocardiogram, and Byar Prostate Cancer with 92%, 78%, and 62% continuous features, respectively), other methods such as FAMD or K-Prototypes performed equally well or better.
Even in cases where DIBmix was not the top method, it produced non-zero and often competitive ARI values, extracting meaningful cluster structure. For example, in the Inflammation and Credit Approval data sets, it outperformed some but not all competitors.
@@ 6 Conclusion
In this paper, we introduced the Deterministic Information Bottleneck algorithm and employed it to devise a new method for clustering mixed-type data. We have described the mathematical foundations of the proposed method and have described the effect of the hyperparameters involved on the algorithm’s performance. We developed a strategy for selecting the hyperparameter values associated with the algorithm, so as to ensure that the different variable types contribute equally to the partition obtained. Moreover, we have ensured that a valid partition into exactly $C$ clusters emerges by recursively updating the value of the regularisation parameter $\beta$. The method has demonstrated promising results in a series of simulations in comparison to four state-of-the-art clustering algorithms for heterogeneous features. Additionally, the application of the algorithm to publicly available data sets has yielded reasonably good results, establishing it as a strong candidate among the clustering methods for mixed-type data proposed to date. Additionally, the algorithm can be readily adapted for clustering solely continuous or categorical data.
Despite DIBmix’s strong performance relative to state-of-the-art clustering methods for mixed-type data, several important limitations warrant discussion. First, bandwidth selection plays a crucial role in cluster detection, and while our proposed approach yields satisfactory results overall, we observed cases where it failed to recover the underlying cluster structure. This highlights the need for a more sophisticated bandwidth selection process. A second limitation stems from DIBmix’s implicit assumption of cluster homogeneity. This becomes particularly problematic when clusters exhibit significantly different characteristics - for example, when spherical clusters have varying volumes. In such cases, applying uniform bandwidth values across all clusters is suboptimal, as larger clusters may disproportionately influence the bandwidth selection. An ideal solution would adapt bandwidth values based on cluster-specific characteristics such as compactness, allowing the algorithm to better account for variations in cluster structure. Future research could address this limitation by borrowing ideas from nearest-neighbour classification algorithms to incorporate local information [40]. A final limitation of the proposed method is that it requires storing full $n\times n$ similarity matrices. This can be computationally prohibitive as the number of observations $n$ grows. Structured approximations or streaming algorithms that operate with partial distance computations may offer a viable path toward improved scalability.
Nevertheless, DIBmix establishes a flexible framework for extending information-based clustering to heterogeneous data, potentially catalyzing the development of next-generation information-based clustering techniques. One promising direction leverages DIBmix’s bandwidth-based architecture for feature weighting and selection. Specifically, the symmetric properties of the perturbed matrix $\mathbf{P^{\prime}}$ could be exploited for this purpose. The eigengap heuristic from spectral clustering [41] presents a natural approach for identifying feature subsets that generate block-like structures in $\mathbf{P^{\prime}}$. However, this approach faces significant computational challenges, as it would require examining $2^{p}-1$ variable combinations, resulting in prohibitive computational complexity of $\mathscr{O}(2^{p}n^{3})$ or $\mathscr{O}(2^{p}n^{2}p)$ for high-dimensional data $(p\gg n)$. Given these computational constraints and the broader challenges in clustering-based feature selection (see [42, 43] for comprehensive surveys), this remains an important direction for future research.
Robustness presents another promising direction for future work. In this study, equal weights ($p(x)=1/n$) were assigned to all observations, but some data points may require weighting based on their outlyingness. Extending DIBmix to incorporate trimming, akin to the Trimmed K-Means algorithm [44], would involve dynamically updating $p(x)$ and adjusting other probabilistic components at each iteration.
Another area for development is the estimation of the number of clusters using information-theoretic quantities. While the original DIB algorithm employed $\beta$ as a model selection parameter, our adaptation does not support this functionality. However, it is well established that the mutual information term $I(T;Y)$ increases with the number of clusters (see Proposition 1 for a proof and Lemma 1 in B and C, respectively, for additional mathematical properties). Preliminary experiments on synthetic data have shown that if $C^{*}$ represents the true number of well-separated clusters, the rate of increase in $I(T;Y)$ when transitioning from $C^{*}-1$ to $C^{*}$ clusters is greater than the rate observed when transitioning from $C$ to C1C-1 clusters, for all $C\neq C^{*}$. We illustrate the results of our experiments on a hundred randomly generated data sets including four spherical clusters with low overlap (level 0.01) and a hundred data sets with four non-spherical clusters with moderate overlap (level 0.10). Each data set consists of 500 observations, four continuous and four categorical variables, and all four clusters are balanced. DIBmix is run with a hundred random initialisations and the mutual information values $I(T_{C};Y)$ for $C=2$ up to $C=8$ clusters are plotted in Figure 4 in D. In the case of well-separated spherical clusters, the knee of the mutual information curve was detected to correspond to $C=4$ in 97% of the data sets. For elliptical clusters with moderate overlap, the corresponding percentage is 85%. This phenomenon may resonate with readers familiar with the elbow heuristic used in K-Means and other distance-based clustering methods for determining the optimal number of clusters. While more advanced methods for estimating the number of clusters have been proposed for non-model-based clustering (see [45] for a comprehensive survey), their applicability is often limited to continuous data. In contrast, the use of a “knee” heuristic derived from DIBmix could provide a novel approach for developing cluster validation indices that exploit information-theoretic quantities. Such indices could facilitate the estimation of the optimal number of clusters in data sets with mixed-type variables, offering a promising avenue for further research.

### v6

@@ 1 Introduction
The literature on clustering offers several taxonomies of algorithms, categorising them based on key features such as the type of output (e.g., hard or fuzzy clustering) or the methodological approach (e.g., model-based or centroid-based). For mixed-type data, model-based approaches have been widely explored, with the Gaussian-Multinomial mixture model [28] standing out as a prominent example. Non-model-based approaches, often referred to as distance-based methods, provide another robust framework for clustering heterogeneous data. Among these, the K-Prototypes algorithm [18] is particularly notable for combining the strengths of K-Means clustering for continuous variables with the K-Modes method for categorical variables. Ji et al. [20] further enhanced this approach by incorporating attribute weights in the K-Prototypes algorithm through entropy-based feature weighting, demonstrating improved performance particularly on data sets with varying feature importance.
Building on this foundation, Rezaei and Daneshpour [36] proposed a three-fold approach combining density-based initialization, threshold-based similarity measures, and a hybrid assignment strategy. While effective for smaller, feature-rich data sets, their method requires careful threshold tuning and scales poorly to large data sets with few classes. Kar et al. [21] introduced EDMD, which uses Boltzmann’s entropy to compute dissimilarities for nominal and ordinal attributes separately, though its effectiveness depends on sufficient feature diversity. Mousavi and Sehhati [31] proposed GUDMM, employing Jensen-Shannon divergence and mutual information to capture inter-attribute dependencies, but the reliance on pairwise calculations limits scalability for high-dimensional data.
@@
In recent years, deep learning approaches such as Deep Embedded Clustering (DEC) [48] and autoencoder-based methods [30] have emerged as alternatives for clustering. However, their application to mixed-type data presents challenges. Categorical variables require one-hot encoding, which increases dimensionality and may obscure cluster structures. Additionally, numerous hyperparameters make fair comparisons difficult, and these methods often lack interpretability. Given our focus on a theoretically grounded approach that natively handles continuous, nominal, and ordinal variables, we restrict comparisons to established methods with well-understood properties.
@@
The remainder of this paper is structured as follows: Section 2 provides background on the Information Bottleneck method and its deterministic variant. Section 3 presents DIBmix, our proposed extension for mixed-type data, detailing its theoretical framework and algorithmic implementation. Section 4 outlines the selection process of hyperparameter values and Section 5 discusses the simulations performed on artificial data to benchmark the proposed method against other established clustering techniques. In Section 6, we apply the DIBmix method to publicly available data sets and analyse its performance. Section 7 discusses the estimation of the number of clusters using information-theoretic quantities. Finally, Section 8 summarises our findings and suggests avenues for future research.
2 The Information Bottleneck Method
@@
$q^{*}(t\mid x)=\operatornamewithlimits{argmin}\limits\limits_{q(t\mid x)}H(T)-\beta I(Y;T)$ (1)
@@ 3 Proposed Methodology
$\quad\times\prod\limits_{j=p_{c}+p_{u}+1}^{p}K_{o}\left(x_{q,j}=x_{r,j};\nu_{j}\right),\quad q,r=1,\ldots,n,$
@@
1exCompute perturbed similarity matrix $\mathbf{P^{\prime}}=\mathbf{P_{Y\mid X}}$ using generalised product kernels with the bandwidths $\boldsymbol{\theta}=(\boldsymbol{s}^{\intercal},\boldsymbol{\lambda}^{\intercal},\boldsymbol{\nu}^{\intercal})^{\intercal}$.
@@
while $m\leq m^{\text{max}}$ and $\mathrm{converged}\neq 1$ do
@@
  1exif $\mathbf{Q^{(m)}_{T\mid X}}=\mathbf{Q^{(m-1)}_{T\mid X}}$ then
@@
     1exbreak
 $\mathbf{q^{(m)}_{T}}\leftarrow 1/n\times\mathbf{Q^{(m)}_{T\mid X}}\boldsymbol{1}_{n}$
@@
For large datasets, constructing the full similarity matrix $\mathbf{P^{\prime}}$ becomes computationally prohibitive as the associated cost is quadratic in the number of observations $(\mathscr{O}(n^{2}))$. For such cases, we recommend utilising the Nyström approximation, which provides a reconstruction of $\mathbf{P^{\prime}}$ using a subset of $m\ll n$ randomly selected landmark points [47]. This replaces dense matrix operations with low-rank approximations by assuming the factorisation $\mathbf{P^{\prime}}\approx\mathbf{C}\mathbf{W}\mathbf{C}^{\intercal}$, where $\mathbf{C}$ and $\mathbf{W}$ denote the $n\times m$ matrix of similarities between all observations and the landmarks and the $m\times m$ similarity matrix of the landmarks, respectively. Theoretical results indicate that the effective rank of kernel matrices grows slowly; thus, choosing $m\approx\sqrt{n}$ landmarks is sufficient to preserve the relevant spectral properties of the data, while reducing the computational complexity to $\mathscr{O}(n\sqrt{n})$ [5]. The cluster recovery performance differs negligibly for different values of $m$, while lower values render the problem much more feasible in terms of computational cost; see G.
@@ 4.1 Bandwidth Selection
For continuous variables, we determine a reasonable bandwidth value $s$ (assuming $\boldsymbol{s}=s(1,\ldots,1)^{\intercal}$) by seeking a high degree of local smoothing. This is quantified using the average nearest-neighbour kernel ratio $\zeta(\boldsymbol{s})$, defined similarly to Expression (7):
@@
In order to overcome the need for constructing the full perturbed similarity matrix $\mathbf{P^{\prime}}$ for large data sets, a subsampling procedure can be used for calculating $\chi(\boldsymbol{s})$ and $\zeta(\boldsymbol{s})$. More precisely, given a data subsample $b$ of size $\tilde{n}\ll n$, the quantities $\chi_{b}(\boldsymbol{s})$ and $\zeta_{b}(\boldsymbol{s})$ are computed for $b=1,\ldots,B$ and their median values across all subsamples are taken as our estimates for $\chi(\boldsymbol{s})$ and $\zeta(\boldsymbol{s})$. The use of the median is motivated by the fact that extreme values may be encountered for highly non-representative subsamples.
@@ 5 Simulations on Artificial Data
Data generation was performed using the MixSim function from the homonymous package [29], with each scenario replicated 50 times. For continuous data, cluster overlap is defined as in [26]. To incorporate nominal features, quartile-based discretisation was applied. In total, 28,800 data sets were generated. Cluster recovery was assessed using the Adjusted Rand Index (ARI) [19] and the Adjusted Mutual Information (AMI) [46]. We only report the ARI values here for brevity; the same conclusions are drawn by considering the AMI values, which are available in the GitHub link provided earlier.
Figure 2: Violin/box plots of ARI values by method. The dashed horizontal line indicates the median ARI of DIBmix ($\approx$0.340).
@@
Figure 2 illustrates the distribution of ARI values across the five methods. The median ARI is highest for DIBmix (0.340), followed by KAMILA (0.328), K-Prototypes (0.326), and FAMD with K-Means (0.308), suggesting that these methods are more consistent in producing high-quality cluster partitions. Conversely, PAM with Gower’s dissimilarity exhibits a wider spread of ARI values, including several low outliers, indicating less stable performance. A repeated measures analysis of variance in which the four competing algorithms (KAMILA, K-Prototypes, FAMD with K-Means, and PAM/Gower) are compared individually with DIBmix also verifies the improvement in cluster recovery performance with the latter. The partial $\eta^{2}$ values are summarised in Table 1. These effect sizes range from 0.14 to 0.26, all exceeding or substantially exceeding Cohen’s rule-of-thumb threshold of 0.14 for large effects [7]. This indicates that DIBmix consistently achieves meaningfully superior cluster recovery performance compared to each competing method across the diverse simulation scenarios. Notice that $p$-values are not reported as these are almost zero due to the large number of comparisons.
@@
Figure 3 provides a detailed comparison of mean ARI values across varying experimental conditions. Overall, DIBmix outperforms the other methods in a majority of scenarios. However, its performance is relatively less competitive when the proportion of categorical variables is high (see Figure 3(c)). This may be attributable to the bandwidth selection strategy outlined in subsection 4.1; alternative bandwidth choices could potentially enhance performance in these cases. Despite the concerns related to the bandwidth selection for data sets with a large proportion of categorical features, DIBmix still performs comparably to other methods in high-categorical scenarios, where most methods, except FAMD/K-Means, struggle. The dimensionality reduction step in FAMD likely provides an advantage in handling data sets dominated by categorical features. Another noteworthy observation is DIBmix’s robustness to cluster imbalances, as illustrated in Figure 3(f). Unlike KAMILA, K-Prototypes, and FAMD/K-Means, DIBmix effectively handles data sets with unbalanced cluster sizes by allowing partitions to be highly imbalanced and not necessarily favouring equally-sized groups. This robustness arises from the entropy term in its objective function, which minimises for imbalanced clusters, thus allowing exploration of diverse partition structures.
@@ 6 Applications to Publicly Available Data
While artificial data provides controlled settings, it does not cover the full range of scenarios encountered in practice. We therefore complement those experiments with evaluations on publicly available data sets from the UCI Machine Learning Repository [10]. These data sets were originally designed for classification tasks, providing a benchmark for assessing clustering accuracy. The ARI values, averaged over one hundred random initialisations, are summarised in Table 2, with the best result for each dataset shown in bold. Note that Adult/Census Income was the only data set consisting of over a thousand observations, so a Nyström approximation with $m=\lceil\sqrt{n}\rceil=174$ random landmark points and a subsampling procedure with a hundred subsamples of a thousand observations for the choice of hyperparameter values was used. Minimal differences in clustering performance were found for different values of $m$, while the recommended value of $\lceil\sqrt{n}\rceil$ required a reasonable runtime; see G.
@@
Adult/Census Income 0.1749 0.1514 0.0864 0.1558 0.1007
@@
    Best performance occurs when data sets contain a balanced mix of variable types. For instance, in the Heart Disease, Adult, Statlog, and Australian Institute of Sport data sets, which all feature a substantial but not overwhelming proportion of categorical variables, DIBmix outperformed its competitors.
@@
    Performance decreases at the extremes. When data sets are dominated by categorical variables (e.g., Dermatology, with nearly 97% categorical features), or continuous (e.g., Australian Institute of Sport, Echocardiogram, and Byar Prostate Cancer with 92%, 78%, and 62% continuous features, respectively), other methods such as FAMD or K-Prototypes performed equally well or better.
Even in cases where DIBmix was not the top method, it produced non-zero and often competitive ARI values, extracting meaningful cluster structure. For example, in the Inflammation and Credit Approval data sets, it outperformed some but not all competitors. We performed a Friedman test on the results for the ten publicly available datasets at 95% significance level. The null hypothesis of equal ranks is not rejected but the critical difference (CD = 1.929; see Figure 4) indicates that methods must differ by approximately two full rank positions to achieve statistical significance. Even the largest observed difference (DIBmix at rank 2.45 vs. Gower/PAM at rank 3.30) represents less than half this threshold. With only ten datasets spanning diverse applications, sample sizes ($n=61$ to $30162$), dimensionalities ($p=6$ to $34$), cluster structures ($C=2$ to $6$), and proportions of categorical variables (8.338.33% to 97.0697.06%), achieving statistical significance would require one method to dominate nearly all comparisons. Despite this, DIBmix demonstrates practical advantages: best average rank (2.45), top performance on four datasets, and robust results with no near-zero ARI values (minimum ARI = 0.1739). This consistency establishes DIBmix as a competitive clustering method for mixed-type data.
Figure 4: Critical Difference (CD) diagram for five clustering methods on 10 publicly available data sets. Average ranks are shown with lower values indicating better performance. The critical difference (CD = 1.929) indicates the minimum rank difference for significance at $\alpha=0.05$.
@@
It is important to stress that despite its poor performance on certain data sets in terms of ARI, the clustering solution need not be assessed solely on its classification performance. While it is common for classification data sets to be used for benchmarking clustering techniques, the performance of a clustering algorithm is mainly assessed by interpreting its clusters [17]. In fact, the partition that is defined by the labels of a classification data set is usually less optimal than the one obtained using a clustering algorithm, with respect to some desirable properties that a cluster has to possess [6].
A desirable property of DIBmix is that it addresses this interpretive need through variable importance quantification. While bandwidth selection is done so as to ensure that no variable dominates the clustering solution a priori, the proposed algorithm learns a partition in which certain variables naturally emerge as more informative for distinguishing between clusters. Variable importance can be assessed by computing the mutual information between the clustering output and each individual feature, the distribution of the latter being estimated using the input kernel function and the bandwidth value that was determined by DIBmix. For instance, Figure 5 shows that clinically meaningful features like bone metastases, the post-trial survival status of patients, and a prostate-specific biomarker (serum prostatic acid phosphatase) drove the obtained partition, while age or blood pressure measurements contributed minimally.
The runtime analysis provided in F indicates that DIBmix executes efficiently and is sometimes even quicker than its competitors on data sets with few observations. For large data sets, such as the Adult/Census Income, the Nyström approximation is essential; without it, DIBmix needed over eleven hours on this data set. Finally, while the hyperparameter search increases computational cost, the computational time remains within practical limits.
Figure 5: Post-analysis of the Byar prostate cancer data set. Variable importance is quantified by the mutual information between each feature and the partition obtained by DIBmix.
7 Selecting the Number of Clusters
Estimating the number of clusters is a notoriously difficult problem in cluster analysis. Here, we present information-theoretic implementations of two common approaches: the elbow method and the Gap statistic [43].
The mutual information $I(T;Y)$ increases with the number of clusters (see Proposition 1 and Lemma 1 in B and C). Experiments on synthetic data show that the rate of increase is greatest when transitioning from $C^{*}-1$ to $C^{*}$ clusters, where $C^{*}$ is the true number of clusters. We tested this on 100 data sets with four well-separated spherical clusters and 100 with moderately overlapping non-spherical clusters, each containing 500 observations and eight variables (four continuous, four categorical). The knee of the mutual information curve correctly identified $C=4$ in 97% of well-separated cases and 85% of overlapping cases (Figure 6, D).
For the Gap statistic, we compute $\text{Gap}(C)=I(T_{C};Y)-\sum_{b=1}^{B}I(T_{C};Y_{b})/B$ using $B=100$ reference data sets generated by independent variable permutation. Selecting $C$ based on maximum Gap yields the correct number of clusters in 80% of well-separated cases but only 25% with moderate overlap, as the Gap curve plateaus due to the monotonic increase of $I(T;Y)$. The ‘one standard error’ rule improves this to 84% and 34%, respectively (Figure 7, E).
Overall, the knee heuristic outperforms the Gap statistic for DIBmix and offers a promising direction for cluster validation in mixed-type data, where few suitable methods currently exist [4].
@@ 8 Conclusion
In this paper, we introduced DIBmix, a clustering method for mixed-type data based on the Deterministic Information Bottleneck framework. We described its mathematical foundations, analysed the role of its hyperparameters, and proposed a strategy to select them so that different variable types contribute equally to the final partition. By recursively updating the regularisation parameter $\beta$, the method guarantees a valid partition into exactly $C$ clusters. Simulation studies and applications to publicly available datasets show that DIBmix performs competitively with state-of-the-art clustering methods for heterogeneous data. The framework is flexible and can be extended to purely continuous or categorical data, as well as to other data types such as counts through appropriate kernel smoothing techniques [23].
Despite its strong performance, DIBmix has several limitations. Bandwidth selection is critical for detecting cluster structure, and while the proposed approach performs well overall, it may fail in certain scenarios. In addition, the method implicitly assumes cluster homogeneity, which can be problematic when clusters differ substantially in size or compactness, making the use of a single bandwidth suboptimal. Future work could address this issue by incorporating local or cluster-specific information, for example by borrowing ideas from nearest-neighbour methods [15].
Nevertheless, DIBmix provides a flexible framework for information-based clustering of heterogeneous data and opens several directions for future research. These include feature weighting and selection using the spectral properties of the perturbed matrix $\mathbf{P^{\prime}}$, although the associated computational costs remain prohibitive in high-dimensional settings. Another promising direction concerns robustness; extending DIBmix to allow observation-specific weights, for instance through trimming strategies similar to Trimmed K-Means [9], could improve performance in the presence of outliers.